# Asymptotic behavior of the ground state energy of a fermionic Fröhlich multipolaron in the strong coupling limit

### v2

@@ Abstract
In this article, we investigate the asymptotic behavior of the ground state energy of the Fröhlich Hamiltonian for a fermionic multipolaron in the strong coupling limit. We prove that it is given, to leading order, by the ground state energy of the Pekar-Tomasevich functional with fermionic statistics, a much simpler model. Our analysis builds upon [Wel15], which itself used and generalized methods developed in [LT97], [FLST11] and [GW13]. Our main new contribution is two-fold. First, we take into account the fermionic statistics of the multipolaron, by employing a localization method used in [LL05]. Second, we relax an assumption on the external electric and magnetic fields, which is not easily verifiable, unless the fields are periodic. Instead, we allow for general fields that only ensure self-adjointness of the Fröhlich Hamiltonian.
@@ Acknowledgments
The first author (IA) is grateful to Marcel Griesemer for suggesting to work on the strong coupling limit of polarons, to Jeremy Faupin for discussions on non-relativistic quantum electrodynamics that turned out to be inspiring for this work, and to Marcel Griesemer, Robert Seiringer, David Wellig, Benjamin Landon and Andreas Wünsch for numerous discussions on polarons and the strong coupling limit. Both authors acknowledge gratefully Dirk Hundertmark and Semjon Vugalter for numerous discussions and references which helped in completing the arguments of the proofs. The research of (IA) was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project-id 258734477 - SFB 1173. This material is based upon work supported by the National Science Foundation under Grant No. DMS-2424139, while (MH) was in residence at the Simons Laufer Mathematical Sciences Institute in Berkeley, California, during the Fall 2025 semester.
@@ 1.1 Background
A polar crystal consists of alternating positive and negative ions, see fig. 1(a). Here we use the expression ’ion’ in a loose sense, in order to also include covalent semi-conductors, such as GaAs, in addition to metals. An example of a polar crystal is NaCl, also known as table salt.
Now consider an electron traveling through a polar crystal. We assume that the ions in the crystal are not rigidly fixed. As the electron moves along the crystal, it distorts the crystal in its direct environment, see fig. 1(b). The ensemble consisting of the traveling electron and the induced distortions of the lattice is called a polaron. In the case of two or more electrons traveling through a polar crystal, we analogously call the resulting ensemble a bipolaron or a multipolaron, respectively. We describe the lattice vibrations by a (massless) bosonic quantum field. The excitations of the bosonic quantum field are referred to as phonons. In particular, electrons interact with phonons as quantum fields.
@@
It is energetically more favorable for the electrons to deform the lattice in a small region, causing various electrons to stay close together. This results in an attractive force mediated by lattice distortions that operates between the electrons and is counteracted by their Coulomb repulsion. As a consequence, even though electrons repel each other, polarons may attract each other. In order to understand the question of whether polarons attract each other or not, one needs an insight into the behavior of the ground state energies of polarons and multipolarons. A well-known and widely accepted polaron model is the Fröhlich model, derived by Herbert Fröhlich [Frö54] in 1954, see also [Cal13]. In the same year, Pekar [Pek54] argued that, in the strong-coupling limit and in the bipolaron case, the model can be reduced by integrating out the phonon modes, leading to the Pekar-Tomasevich model [PT51, LLP53]. Feynman [Fey55, Fey98] used the Feynman-Kac formula to compute an expansion of the ground-state energy of the Fröhlich model in the coupling constant, and identified the leading order term as the Pekar-Tomasevich energy.
In 1980, Adamowski, Gerlach and Leschke [AGL80] sketched a proof of the convergence of the Fröhlich polaron ground state energy to the Pekar-Tomasevich polaron groundstate energy with the help of large deviation theory of Brownian motion. In 1983, Donsker and Varadhan [DV83] provided a complete and rigorous proof. However, information on the rate of convergence was not provided. In 1997, Lieb and Thomas found in [LT97] a simpler proof which also gives rates of convergence. Miyao and Spohn [MS07] extended the result to the bipolaron case. In 2013, Wellig and Griesemer [GW13] included a polaron with external electric and magnetic fields. In the same year, the multipolaron case with no external fields was studied by the first author and Landon in [AL13]. They combined the methods of Lieb and Thomas and also applied Feynman-Kac formulas. They got inspired by application of the Feynman-Kac formulas in [FLST11] to show that bipolarons do not form a bound state above a threshold of $\nu$. A few months later, Wellig [Wel15] generalized their phonon localization method to multipolarons, and included the external field dependence, by adapting the single polaron analysis [GW13] to the multipolaron setting.
Neither [AL13] nor [Wel15] took the fermionic statistics into account. Furthermore, [Wel15] relied on an an energy assumption, see (1.24) below, which is not easy to verify unless the external fields are constant or periodic. Our contribution is on the one hand to take the fermionic statistics of the electrons into account and on the other hand to remove the energy assumption. We have to pay with the price that the error terms grow faster in $N$, namely as $N^{\frac{82}{30}}$. The error presented in [Wel15] grows as $N^{3}$; however, optimizing parameters there leads to an error growing as $N^{\frac{57}{23}}$.
@@
Ghanta [Gha15] proved that in the one-dimensional case for a single electron ($N=1$), in the absence of an external magnetic field ($A=0$), and some assumptions on $V$, approximate minimizers of the Fröhlich Hamiltonian converge to the, up to a rotation and translation unique, minimizer of the Pekar-Tomasevich functional in some appropriate topology. In this work, we do not consider this convergence.
Another direction, which recently has attracted a lot of attention, lies in the study the effective mass of the Fröhlich polaron, which is increased compared to the electron mass without the coupling, [BMSV25, Bro24, Sel24], and the related study of the energy dependence on the total momentum [BS24, LMM23, MMS23, Pol23]. Brooks and Seiringer [BS23] have recently studied the associated Bogoliubov theory. For an asymptotic low-energy expansion, see [BM25, Sei21]. We do no pursue any of these directions in this work, but it would be interesting extending our ideas to these different problems.
1.2 Polarons in the continuum limit: The Fröhlich model
Lattice vibrations associated with out-of-phase movements of neighboring atoms, e.g., in a two-atomic crystal are referred to as optical phonons, see fig. 2(a). Moreover, in 3 dimensions, the phonon momentum can be aligned or orthogonal to the electron momentum. We are interested in the case that electrons are coupled to optical phonons with aligned momentum, referred to as longitudinal optical phonons. Near the origin, their dispersion is nearly constant, see fig. 2(b) for the 1 dimensional case. The physical value is referred to as Debye frequency, which we set to 1 in our units. We present the ideas in the continuum limit, for which the lattice constant goes to zero. However, it would be possible to repeat our analysis in the discrete setting, as originally presented in [LT97]. The resulting model describing electrons coupled to longitudinally polarized optical phonons in a polar crystal is referred to as the Fröhlich model [Frö54].
@@ 1.2 Polarons in the continuum limit: The Fröhlich model
Note that the assumption $A_{k}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{3})$, $k=1,2,3$ ensures that $D_{A_{\alpha},x}$ is self-adjoint with domain
$H^{1}_{A_{\alpha}}(\mathbb{R}^{3})\,:=\,\{f\in L^{2}(\mathbb{R}^{3})\mid\forall k=1,2,3:\,(-i\partial_{x_{k}}+(A_{\alpha})_{k})f\in L^{2}(\mathbb{R}^{3})\}\,$, (1.6)
see [LL96, Section 7.20-7.22].
Electrons are described by the fermionic Hilbert space $\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})$ of wavefunctions $\Psi$ satisfying almost everywhere the equality
$\Psi(x_{\pi(1)},\ldots,x_{\pi(N)})\,=\,\operatorname{sgn}(\pi)\Psi(x_{1},\ldots,x_{N})$ (1.7)
for all permutations $\pi\in\mathcal{S}_{N}$ of labels $\{1,\ldots,N\}$. $a$, $a^{\dagger}$ are the bosonic creation and annihilation operators on the bosonic Fock space
@@
where $\bigvee_{j=1}^{n}L^{2}(\mathbb{R}^{3})$ denotes the symmetric subspace of $L^{2}(\mathbb{R}^{3n})$. Every $\eta_{n}\in\bigvee_{j=1}^{n}L^{2}(\mathbb{R}^{3})$ satisfies almost everywhere the equality
@@
It describes electron scattering by a phonon with momentum transfer $k\in\mathbb{R}^{3}$. More precisely, annihilation of a phonon of momentum $k$ results in an electron gaining momentum $k$. Creation of a phonon of momentum $k$ results in an electron losing momentum $k$. The form factor $(\sqrt{2}\pi|k|)^{-1}$ is tied to the scattering of an electron with momentum $k$ scattering with longitudinally polarized optical phonons. It is a product of the electron momentum $|k|$, the Fourier transform of the Coulomb potential $(2\pi^{2}|k|^{2})^{-1}$, and the phonon amplitude $\pi\sqrt{2\omega(k)}$. The dispersion $\omega$ of longitudinal optical phonons is approximately constant – its value is known as the Debye frequency – and units are set such that $\omega(k)=1$. For more details, we refer to [Fey98, Chapter 8.1].
Since the dispersion is constant, $\omega(k)=1$, the phonon field energy of is given by the number of excitations, namely
@@
The total Hilbert space describing both, phonons and electrons, is given by
@@
A difficulty in realizing $H^{(N,\alpha)}$ as a self-adjoint operator stems from the fact that $\frac{e^{ix\cdot}}{|\cdot|}\notin L^{2}(\mathbb{R}^{3})$. To overcome this obstacle, one casts $H^{(N,\alpha)}$ as a quadratic form $q^{(N,\alpha)}$ on $\mathcal{Q}_{0}$ and proves that it is closable. In fact, $\phi$ is infinitesimally relatively form-bounded w.r.t. $D_{\tilde{A}}^{2}+\mathcal{N}$. Then the KLMN theorem, see [RS75, Theorem X.17, p. 167], [Lew24, Theorem 3.19], yields that $q^{(N,\alpha)}$ has a unique closed extension on $\mathcal{H}_{N}$. This closed extension then is associated with a unique self-adjoint operator that is bounded from below, see [BS87, Theorem 2, Chapter 10]. For more details, we refer to [Hot16, MS07, Nel64]. For a description of the domain of self-adjointness, we refer to [GW16].
By the Courant-Fischer min-max principle, the ground state energy $E^{(N,\alpha)}:=\inf\sigma(H^{(N,\alpha)})$ satisfies
@@
1.3 Statement of main result: Strong-coupling limit
@@
We then denote the ground state energy in the Pekar-Tomasevich model by
@@
Let
denote the density of $\Psi_{N}\in\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})$. A now standard calculation originally found in [Pek54], see Lemma A.1, implies that
Our goal is to show that the Pekar-Tomasevich functional, which is a simpler model than the full Fröhlich Hamiltonian, yields the behavior of the ground state energy of the Fröhlich Hamiltonian up to a relative error which vanishes in the strong coupling limit $\alpha\to\infty$. In view of (1.19), an appropriate lower bound for the ground state energy of the Fröhlich Hamiltonian is sufficient. Such a lower bound is provided in our main Theorem 1.1.
1.1 Theorem.

### v3

# arXiv:1601.05272v3[math-ph] 28 Feb 2026

@@ Abstract
In this article, we prove that the ground-state energy of a fermionic Fröhlich multipolaron can be approximated, in the strong electron-phonon coupling limit, by the ground-state energy of a corresponding fermionic Pekar-Tomasevich multipolaron, even in the presence of external electric and magnetic fields. Our analysis builds upon Lieb and Thomas’ approach [29], which was originally developed for a single polaron without external fields, and Wellig’s generalization to multipolarons [42] with (specialized) external fields. Our main new contributions are twofold. First, we take into account the fermionic statistics of the multipolaron by employing a localization method from [28]. Second, we relax an assumption in [42] on the external electric and magnetic fields, which is not easily verifiable unless the fields are periodic. Instead, we allow for general fields that only ensure self-adjointness of the Fröhlich Hamiltonian. In particular, our work demonstrates the robustness of the Lieb–Thomas strategy when extended to fermionic multipolarons and general external potentials.
@@ Acknowledgments
The first author (IA) is grateful to Marcel Griesemer for suggesting that we work on the strong-coupling limit of polarons, to Jeremy Faupin for discussions on non-relativistic quantum electrodynamics that turned out to be inspiring for this work, and to Marcel Griesemer, Robert Seiringer, David Wellig, Benjamin Landon and Andreas Wünsch for numerous discussions on polarons and the strong-coupling limit. Both authors gratefully acknowledge Dirk Hundertmark and Semjon Vugalter for numerous discussions and references that helped complete the arguments of the proofs. The research of (IA) was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 258734477 – SFB 1173. This material is based on work supported by the National Science Foundation under Grant No. DMS-2424139; (MH) was in residence at the Simons Laufer Mathematical Sciences Institute in Berkeley, California, during the Fall 2025 semester.
@@ 1.1 Background
Consider an electron moving through a polar crystal (Fig. 1(a)). Such a crystal consists of alternating positive and negative ions; a standard example is NaCl (table salt). We assume the ions are not rigidly fixed. As the electron moves, it distorts the lattice in its immediate neighborhood (Fig. 1(b)). The electron together with this induced distortion is called a polaron. For two or more electrons, the corresponding objects are bipolarons or multipolarons. We model the (quantized) lattice vibrations by a massless bosonic field; its excitations are phonons, and electrons interact with them through the field.
@@
It is energetically favorable for electrons to deform the lattice in a localized region, which can keep them close together. The resulting distortion produces an effective attraction that competes with Coulomb repulsion; thus polarons may attract even though bare electrons repel. To study effective multipolaron interactions, we focus on the ground-state energy. A widely accepted model is the Fröhlich model, proposed by Herbert Fröhlich [16] in 1954; see also [10]. In the same year, Pekar [35] argued that in the strong-coupling limit the phonon modes can be integrated out; see also [24]. Together with Tomasevich [34], this yields the effective Pekar-Tomasevich multipolaron model, with the single-polaron case known as the Pekar polaron. Feynman [13, 14] used a Feynman–Kac formula to expand the Fröhlich ground-state energy in the coupling constant and identified the leading term as the Pekar polaron energy.
In 1980, Adamowski, Gerlach and Leschke [1] sketched a proof of convergence of the Fröhlich polaron ground-state energy to the Pekar-Tomasevich energy using large deviations for Brownian motion. In 1983, Donsker and Varadhan [12] gave a complete proof, but without rates. In 1997, Lieb and Thomas [29] provided a simpler argument that also yields convergence rates. Their approach applies Pekar’s idea to a UV-truncated and spatially localized Fröhlich model and controls the truncations in the strong-coupling limit. We adapt their strategy, which proceeds in the following steps.
  1. (i)
    Electron localization: Up to an energy error $\Delta E$, electrons can be localized inside a cube of side-length $\sim(\Delta E)^{-\frac{1}{2}}$, in accordance with Heisenberg’s uncertainty principle.
  2. (ii)
    UV cutoff: Using a commutator estimate, see Lemma 3.1, they truncate UV phonon modes with momentum above $\Lambda$.
  3. (iii)
    Block-mode reduction: To reduce the number of phonon modes, they dissect the $\Lambda$-cube into $(\Lambda/P)^{3}$ blocks of side length $P$; see Prop. 3.2. Using translation invariance of the Hamiltonian, the interaction per block reduces to that of an electron with a single representative block mode in the cube.
  4. (iv)
    Phonon integration: Using coherent state analysis, they then employ Pekar’s argument, which is based on completion of the square, to this reduced Hamiltonian, see Prop. 3.3 and, for a more detailed explanation, Appendix A.
Both the many-particle case and the presence of external potentials introduce new challenges. With many particles, interparticle distances create an additional length scale that can interfere with the localization step. External potentials break translation invariance, which underlies the block reduction. Moreover, the final Lieb–Thomas reduction uses the scale invariance of the Pekar functional, which is violated by both the Coulomb interaction and external fields.
In 2013, Wellig and Griesemer [19] studied the single polaron with external electric and magnetic fields. In the same year, the first author and Landon [3] generalized the Lieb–Thomas argument to the multipolaron case without external fields. To handle many particles, they replaced the electronic localization step with
  1. (i’)
    Cluster localization: Split the multipolaron into various localized clusters.
Inspired by [15], they use Feynman–Kac formulas to bound the energy of interacting clusters by that of non-interacting clusters. To close the Lieb–Thomas argument, they add:
  1. (v)
    Effective collective cluster energy: Bound the sum of Pekar-Tomasevich energies of the clusters by the total energy of the Pekar-Tomasevich multipolaron, see Corollary 3.7.
We denote by $E_{PT}^{(n,\alpha)}$ the ground-state energy of the Pekar–Tomasevich nn–polaron with coupling constant $\alpha$ between the electrons and the phonon field. In the absence of external fields, the Pekar-Tomasevich functional is translationally invariant and consequently obeys the subadditivity
$E_{PT}^{(n,\alpha)}\leq E_{PT}^{(k,\alpha)}+E_{PT}^{(n-k,\alpha)},\quad\forall 1\leq k<n\leq N\,.$ (1.1)
We explain this in more detail in Rem. 3.6. As a corollary of strong-coupling convergence, they show that bipolarons form a bound state if the effective Coulomb coupling $\nu$ is below a threshold $\nu_{c}>2$; see (1.3).
A few months later, Wellig [42] extended [3] to include external potentials. A key difficulty is that the Feynman–Kac formula is no longer available, and (1.1) need not hold. In the field-free case, Feynman–Kac allows simultaneous localization of electronic support and interacting phonon modes. Instead, Wellig first proves a covering lemma (Lemma 2.1) and develops a cluster localization argument similar to [3] (see Prop. 2.2). He then generalizes the proof to arbitrary clusters by localizing phonons to regions near the electron clusters they interact with, following [15]; see Lemma 3.2. To carry out step v, he assumes (1.1) for the external fields; periodic fields provide a notable example.
Note that neither [3] nor [42] allow for fermionic statistics, since their localization functions do not preserve particle statistics. We overcome this challenge by implementing localization functions used in [28] in the cluster localization argument i, at the cost of larger error terms. Our error is of order $O(N^{\frac{82}{30}})$ as $N\to\infty$, while the error presented in [42] grows as $O(N^{3})$; however, optimizing parameters there leads to an $O(N^{\frac{57}{23}})$ error. To remove (1.1), we exploit the fact that cluster localization can be preserved throughout the phonon integration, i.e., in the reduced Pekar-Tomasevich model, in the Lieb–Thomas argument. Since the clusters are well separated, their interaction can be controlled, allowing us to replace (1.1) with the estimate
$E_{PT}^{(N,\alpha)}\,\leq\,\sum_{j=1}^{m}E_{PT}^{(n_{j},\alpha)}(B_{j})+C\frac{N^{2}\alpha}{R},\quad\sum_{j=1}^{m}n_{j}=N\,,\quad\min_{j\neq k}\operatorname{dist}(B_{j},B_{k})\geq R\,$, (1.2)
see Corollary 3.7. Here $E_{PT}(B_{j})$ denotes the Pekar-Tomasevich ground-state for multipolarons confined to a ball $B_{j}\subseteq\mathbb{R}^{3}$. This estimate holds for very general external potentials and imposes no additional constraints beyond self-adjointness of the Fröhlich Hamiltonian.
@@
Miyao and Spohn [31] proved strong resolvent convergence of the UV cut-off Fröhlich Hamiltonian to the full model and studied existence of the ground-state energy at fixed total momentum. Ghanta [17] showed that in one dimension with a single electron ($N=1$), no magnetic field ($A=0$), and suitable assumptions on $V$, approximate minimizers of the Fröhlich Hamiltonian converge (up to rotation and translation) to the unique minimizer of the Pekar-Tomasevich functional in an appropriate topology. We do not address this convergence here.
Another active direction concerns the effective mass of the Fröhlich polaron, which increases relative to the uncoupled electron mass [4, 6, 39], and the related dependence of the energy on total momentum [9, 23, 30, 36]. Brooks and Seiringer [8] studied the associated Bogoliubov theory. For asymptotic low-energy expansions, see [7, 38]. We do not pursue these directions here, but extending our ideas to them would be interesting.
1.2 Definition of models
Lattice vibrations with neighboring atoms moving out of phase (e.g., in a diatomic crystal) are called optical phonons; see Fig. 2(a). In three dimensions, phonon momentum can be aligned with or orthogonal to the electron momentum. We focus on the aligned case, i.e., longitudinal optical phonons. Near the origin, the dispersion is nearly constant; see Fig. 2(b) for the one-dimensional case. This constant is the Debye frequency, which we set to 1. We work in the continuum limit (lattice spacing $\to 0$), though the analysis can be repeated in a discrete setting as in [29]. The resulting model of electrons coupled to longitudinal optical phonons in a polar crystal is the Fröhlich model [16].
@@ 1.2 Definition of models
The assumption $A_{k}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{3})$, $k=1,2,3$ ensures that $D_{A_{\alpha},x}$ with domain
$H^{1}_{A_{\alpha}}(\mathbb{R}^{3})\,:=\,\{f\in L^{2}(\mathbb{R}^{3})\mid\forall k=1,2,3:\,(-i\partial_{x_{k}}+(A_{\alpha})_{k})f\in L^{2}(\mathbb{R}^{3})\}$ (1.8)
is self-adjoint, see [27, Section 7.20-7.22].
Electrons are described by the fermionic Hilbert space $\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})$ of wavefunctions $\Psi$ satisfying for almost all $x_{1},\ldots,x_{N}\in\mathbb{R}^{3}$ and all permutations $\pi\in\mathcal{S}_{N}$ of labels $\{1,\ldots,N\}$ the equality
$\Psi(x_{\pi(1)},\ldots,x_{\pi(N)})\,=\,\operatorname{sgn}(\pi)\Psi(x_{1},\ldots,x_{N})\,.$ (1.9)
$a$, $a^{\dagger}$ are the bosonic creation and annihilation operators on the bosonic Fock space
@@
where $\bigvee_{j=1}^{n}L^{2}(\mathbb{R}^{3})$ denotes the symmetric subspace of $L^{2}(\mathbb{R}^{3n})$. An element $\eta_{n}\in\bigvee_{j=1}^{n}L^{2}(\mathbb{R}^{3})$ satisfies almost everywhere the equality
@@
It describes electron scattering by a phonon with momentum transfer $k\in\mathbb{R}^{3}$. More precisely, annihilation of a phonon of momentum $k$ results in an electron gaining momentum $k$. Creation of a phonon of momentum $k$ results in an electron losing momentum $k$. The form factor $(\sqrt{2}\pi|k|)^{-1}$ is tied to the scattering of an electron with momentum $k$ by longitudinally polarized optical phonons. It is a product of the electron momentum $|k|$, the Fourier transform of the Coulomb potential $(2\pi^{2}|k|^{2})^{-1}$, and the phonon amplitude $\pi\sqrt{2\omega(k)}$. The dispersion $\omega$ of longitudinal optical phonons is approximately constant – its value is known as the Debye frequency – and units are set such that $\omega(k)=1$. For more details, we refer to [14, Chapter 8.1].
Since the dispersion $\omega(k)=1$ is constant, the phonon field energy is given by the number of excitations, namely
@@
The total Hilbert space describing both phonons and electrons is given by
@@
A difficulty in realizing $H^{(N,\alpha)}$ as a self-adjoint operator stems from the fact that $\frac{e^{ix\cdot}}{|\cdot|}\notin L^{2}(\mathbb{R}^{3})$. To overcome this obstacle, one casts $H^{(N,\alpha)}$ as a quadratic form $q^{(N,\alpha)}$ on $\mathcal{Q}_{0}$ and proves that it is closable. In fact, $\phi$ is infinitesimally relatively form-bounded w.r.t. $D_{\tilde{A}}^{2}+\mathcal{N}$. Then the KLMN theorem, see [37, Theorem X.17, p. 167], [26, Theorem 3.19], yields that $q^{(N,\alpha)}$ has a unique closed extension on $\mathcal{H}_{N}$. This closed extension then is associated with a unique self-adjoint operator that is bounded from below, see [5, Theorem 2, Chapter 10]. For more details, we refer to [21, 31, 33]. For a description of the domain of self-adjointness, we refer to [20].
By the Courant-Fischer min-max principle, the ground-state energy $E^{(N,\alpha)}:=\inf\sigma(H^{(N,\alpha)})$ satisfies
@@
Let
denote the density of $\Psi_{N}\in\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})$. A now standard calculation originally found in [35], see Lemma A.1, implies that
We then denote the ground-state energy in the Pekar-Tomasevich model by
@@
1.1 Remark (Homogeneity of $E_{PT}^{(N,\cdot)}$).
$\mathcal{E}_{PT}^{(N,\alpha)}(\Psi_{\alpha})\,=\,\alpha^{2}\mathcal{E}_{PT}^{(N,1)}(\Psi)\,.$ (1.25)
Consequently, the Pekar-Tomasevich energy satisfies $E_{PT}^{(N,\alpha)}=\alpha^{2}E_{PT}^{(N,1)}$.
1.3 Main result
Our goal is to show that the Pekar-Tomasevich functional, which is a simpler model than the full Fröhlich Hamiltonian, yields the behavior of the ground-state energy of the Fröhlich Hamiltonian up to a relative error which vanishes in the strong-coupling limit $\alpha\to\infty$. In view of (1.24), an appropriate lower bound for the ground-state energy of the Fröhlich Hamiltonian is sufficient. Such a lower bound is provided in our main result, Theorem 1.2.
1.2 Theorem (Validity of Pekar-Tomasevich reduction in strong-coupling limit).